OpenStudy (anonymous):

What is the inverse of g(x) = x^2-5x-14 Help please :)

OpenStudy (helder_edwin):

u mean its inverse function?

OpenStudy (anonymous):

it is not a one to one function, it does not have an inverse however you can still solve \[x=y^2+5y-14\] for \(y\) if you like

OpenStudy (anonymous):

I have to restrict the domain and range so I can find the inverse....doing my head in LOL

OpenStudy (helder_edwin):

ok.

OpenStudy (helder_edwin):

so u have \[\large y=x^2-5x-14 \] \[\large y+14+25/4=x^2-5x+\frac{25}{4} \] \[\large y+81/4=(x-5/2)^2 \]

OpenStudy (helder_edwin):

this shows that the function is symetric in respect with x=5/2. therefore u can restrict the domain to either \([5/2;\infty)\) or \((-\infty;5/2]\).

OpenStudy (helder_edwin):

sorry *with respect to

OpenStudy (anonymous):

where does 25/4 come from?

OpenStudy (helder_edwin):

from completing the square

OpenStudy (helder_edwin):

u still have to solve for \(x\) the equation \[\large y=f(x)=x^2-5x-14 \]

OpenStudy (anonymous):

so the inverse is y+81/4=(x-5/2)^2?

OpenStudy (helder_edwin):

no. that was just some calculations to find out the restriction of the domain. u have to solve the equation \[\large y=f(x)=x^2-5x-14 \] for \(x\).

OpenStudy (anonymous):

x is 7 or -2

OpenStudy (helder_edwin):

no. u have basicly a quadratic equation for x: \[\large x^2-5x-(14+y)=0 \] using the quadratic formula u get \[\large x=\frac{5\pm\sqrt{(-5)^2-4\cdot1\cdot(-(14+y))}}{2\cdot1} \]

OpenStudy (helder_edwin):

\[\large x=\frac{5\pm\sqrt{81+4y}}{2} \]

OpenStudy (helder_edwin):

so the inverse function would be either \[\large f^{-1}(y)=\frac{5+\sqrt{81+4y}}{2} \] or \[\large f^{-1}(y)=\frac{5-\sqrt{81+4y}}{2} \] deppending on which interval u chose for the domain.

OpenStudy (anonymous):

x\[x \ge-\frac{ 81 }{ 4 } y \ge \frac{ 5 }{ 2 }\]

OpenStudy (helder_edwin):

domain corresponds to \(x\) and the range to \(y\).

OpenStudy (anonymous):

so which inverse function would I choose?

OpenStudy (helder_edwin):

if u choose \([5/2;\infty)\) for the domain then \[\large f^{-1}(x)=\frac{5+\sqrt{81+4x}}{2} \]

OpenStudy (anonymous):

Thank you so very very much for all of your help, I really appreciate it :)

OpenStudy (helder_edwin):

u r welcome

OpenStudy (anonymous):

so I just sub in 5/2 in the equation and see what I get as to whether it satisfies the domain ?

OpenStudy (helder_edwin):

NO. the domain is the set of points (real numbers) for which the equation (function) makes sense. since your function has no restrictions (logarithms, divisions, or roots of even power) its domain is actually the whole of \(\mathbb{R}\). BUT. since the function is not one-to-one on this domain, u were told to "reduce" its domain to a set upon which the function is one-to-one.